OpenStudy (anonymous):

If f(x)=6x - 8, g(x)=1/2x, and h(x)=x+4, find (f o g o h)(x). My answer is 3x^3 + 8x^2 - 16x. Is it correct?

sam (.sam.):

f(x)=6x - 8, g(x)=x/2, and h(x)=x+4, find fgh(x)??

OpenStudy (anonymous):

yes

sam (.sam.):

\[fg(x+4)\] \[f(\frac{1}{2}(x+4))\] \[f(\frac{x+4}{2})\] \[6(\frac{x+4}{2})-8\] 3x+12-8 3x+4